OpenStudy (anonymous):

What is the equation of the parabola, in vertex form, with vertex at (2,-4) and directrix y = -6?

OpenStudy (tkhunny):

Have you considered the definition? \(\sqrt{(x-2)^2+(y+4)^2} = y + 6\)

OpenStudy (anonymous):

no however i still dont understand.

OpenStudy (anonymous):

hello @tkhunny

OpenStudy (tkhunny):

Do the algebra to get that expression onto the standard form. What I gave is the focus-directrix definition of a parabola. The left hand side is the distance to the focus. The right hand side is the distance to the directrix. For a parabola, these two are equal. Square both sides and simplify. You'll see it.

OpenStudy (anonymous):

i have no idea how to solve this still @tkhunny

OpenStudy (tkhunny):

Did you square both sides? Everything is positive. It will be okay.

OpenStudy (anonymous):

so id get (x+2)+(y+4)=(y+6)^2?? @tkhunny

OpenStudy (tkhunny):

Not quite. Where did the exponents go - the ones that used to be under the radical? You should have: \( (x-2)^{2}+(y+4)^{2}=(y+6)^2?? \) And why did you change x-2 to x+2? Don't do that.

OpenStudy (anonymous):

oh right sorry.

OpenStudy (anonymous):

ok so would the equation be (x-2)^2 =8(y+4) @tkhunny

OpenStudy (anonymous):

help @tkhunny

OpenStudy (tkhunny):

\((x-2)^{2} = (y+6)^{2} - (y+4)^{2}\) \((x-2)^{2} = y^{2} + 12y + 36 - (y^{2} + 8y + 16)\) \((x-2)^{2} = 12y + 36 - 8y - 16)\) \((x-2)^{2} = 4y + 20)\) \((x-2)^{2} = 4(y + 5))\) This puts the vertex at (2,-5), which is EXACTLY the middle if the distance between the focus and the directrix. Hey, look!! We just invented the standard formula! \((x-h)^{2} = 4p(y + k))\) With h = 2, k = -5, and p = 1

OpenStudy (anonymous):

wow thank you for all your help its been amazing.